OpenStudy (anonymous):

The number of farms in a certain area can be represented by the equation, y = 85,000(0.97)x, where x represents the number of years after 1995. Calculate and show the number of farms in 2020 (x = 25). Does this look like a reasonable number if the initial number in 1995 was 100,000 farms and the number is declining each year? yes or no

OpenStudy (anonymous):

2020 is 25 years after 1995 so your new equation becomes: y=85,000(0.97)25 Solve for y.

OpenStudy (anonymous):

y = 85,000(0.97)x \[\huge y = 85,000(0.97)^x\] ?

OpenStudy (anonymous):

Oh....

OpenStudy (anonymous):

y = 85,000(0.46697470525)